OpenStudy (anonymous):

Using elimination, what is the solution of the system? 4x + 3y = -19 3x - 2y = -10 a) (-4,-1) b) (-1,-4) c) (4,17) d) (17,4)

OpenStudy (anonymous):

have you learned about matrices and linear algebra :)?

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

yes, what do you need about this subject.

jhonyy9 (jhonyy9):

sousing elimination method need multiplie first equation by 2 and second by 3 and after this assum like : 4x +3y =-19 / 2 3x -2y = -10 / 3 8x +6y =-38 9x -6y = -30 ------------ 17x 0 = -68 17x=-68 x= -68/17 hope that this will help

OpenStudy (anonymous):

the solution is (a), but try to think in matrices to solve the problem. Think coeffecient. You could also solve it without matrices and so forth. Many was to do it :)

jhonyy9 (jhonyy9):

there has wrote that need to solve by elimination

OpenStudy (anonymous):

Thank you @jhonyy9 and @Servet!!!

OpenStudy (anonymous):

no probs. @jhonyy9 you are right :)